# Characterizing Continual Learning Scenarios and Strategies for Audio Analysis

### v1

@@ IV-B Non-Continual learning baselines
IV-B1 Joint-training
This is a popular approach in non-CL category, where all the training episodes are jointly trained one time and achieve the highest average performance and hence also regarded as an upper bound. However, this method assumes the availability of all the episodes simultaneously which is a bottleneck.
IV-B2 Cumulative learning
This non-CL training strategy is another widely adopted approach offering upper-bound performance. Here, the training process is started from scratch each time a new episode is available. It considers the current and all the previous episodes together for retraining. However, previously learned knowledge is wasted due to retraining from scratch each time a new episode is encountered leading to high computation requirements.
IV-B3 Naive
This non-CL strategy achieves lower-bound performance. Here, a traditional transfer learning mechanism is followed. Whenever a new episode is available, the model adapts itself to this new data without following any mechanism to control forgetting on past data.
@@ IV-B3 Naive
TABLE II: Details of datasets
@@ V-B Dataset for CI Scenario
Contrary to the domain-incremental dataset discussed in the earlier section, here, the end goal is machine-type classification. In the class-incremental learning scenario, there needs to be the inclusion of classification classes over time, hence we opt to classify machine type as the DCASE datasets offer a large number of machine types which can then be sequentially learned by the CL model. To focus only on machine-type classification, either normal or abnormal data should be considered. Since the normal class data is available in large amounts, hence we consider the normal class samples of different machine sounds here. Specifically, we curate six episodes by utilizing the training part (it has only normal samples) of the development segment in task-2 of DCASE2020 and the source domain of the additional training segment in task-2 of DCASE2023 challenges. The choice of these datasets is such that they have disjoint machine classes. Specifically, we consider ToyCar and ToyConveyer machine types from DCASE2020 in the first episode, Valve and Fan machine types in the second episode, and then Pump and Slider machine types in the third episode. Then from the DCASE2023, two machine types viz., Vacuum and ToyTank are considered in the fourth episode, ToyNscale and ToyDrone are considered in the fifth episode, and finally, the last episode has the remaining classes from DCASE2023, including Bandsaw, Grinder, and Shaker. Inside every episode, the number of samples for each class is considered in balance amount, i.e., the machine types have 3000, 2804, 990, 990, and 990 samples in the first, third, fourth, fifth, and sixth episodes, respectively. Except for the second episode in which the two machine types, namely Valve and Fan, have 3000 and 2804 samples, respectively. A visual representation of the class-incremental dataset is depicted in Fig. 3. The detailed description of datasets of both scenarios is shown in Table II.
VI Experiments
VI-A Experimental setup
@@
In all the approaches, we keep the base model as a pre-trained ResNet50 [58] model initialized with ImageNet weights. We used Adam as an optimizer. In order to select the best hyperparameters for the classification model, multiple experiments were performed with a cumulative approach as it gives the upper bound. Specifically, we search best values for epochs in {10, 15,…50}, batch sizes in {4,8,…32}, and learning rates in {$041e-$ 04, … $051e-$ 05}. Best performance is achieved with a batch size of 8, a learning rate of $041e-$ 04, and epochs of 30 for CI scenario. For DI scenario, the best performance is achieved with a batch size of 8, a learning rate of $031e-$ 03, and epochs of 50. To facilitate a common evaluation scheme the selected hyperparameters are used in all the approaches. Further, the hyperparameters for CL approaches such as $\lambda$, memory size, etc. were also experimentally selected from a range of values. Specifically, for DI experiment, we set $\lambda=0.5$ in EWC, $\alpha=2$, $T=2$} in LwF, and $\lambda=0.8$ in SI. In case of CI experiment, $\lambda=2$ was set in EWC, $\alpha=2$, $T=2$} kept for LwF, and $\lambda=2$ for SI. Further, for both DI and CI, the memory sizes were kept as 2000 samples for GDumb and Replay approaches. Audio clips were trimmed to 10 seconds and then spectrograms were computed on each sample.
@@
VI-B Evaluation metrics
Here, we describe evaluation procedure and metrics in CL. Each of the training session is a classification problem therefore we compute accuracy on the testing data of each task. After completing the training session of $t^{th}$ task using training data $Tr_{t}$, we compute accuracy values on testing datasets $Te_{1},Te_{2},\dots$ yielding $T$ accuracy values. Thus, after completing all the training sessions on a stream of $T$ tasks, we get a train-test matrix $R^{T\times T}$, where $R_{t,j}$ represents accuracy on $j^{th}$ task after completing the training on $t^{th}$ task. An example of train-test matrix with 5 tasks is shown in Fig. 4.
Apart from getting average accuracy across tasks, other important evaluation measures in CL includes forgetting and forward transfer, which can be computed from this $R$ matrix. Below, we describes metrics for forward transfer, forgetting, and model accuracy.
Backward Transfer (BWT): BWT, a forgetting measure, quantifies the amount of influence of the current task on the previously learned task [14]. It shows the improvement in performance of a task after learning news tasks. Thus, a positive value of BWT signifies that learning of current task $t$ has caused improvement in performance of an old task $k<t$. On the other hand, a negative value of BWT indicates that learning of current task $t$ has caused degradation in performance of an old task. Thus, algorithm with greater values of BWT is superior to the one with small BWT. Blue and yellow cell values in Fig. 4 are utilized to compute BWT using the equation [59] below.
@@ VI-B Evaluation metrics
Forward Transfer (FWT): FWT signifies the amount of influence of the current task on upcoming tasks. It shows the zero-shot learning capabilities of the CL model. Higher the value of FWT, better is the model. Red cell values in Fig. 4 are utilized to compute FWT [59] measure with the following equation.
@@
TABLE III: Performance comparison for DI scenario. First / second best performance in CL categories indicated in red / blue , respectively. Bold: Upper bound.
@@
VI-C Results
VI-C1 Performance analysis for DI scenario
For DI scenario, we compute average accuracy over all tasks after completing a training session and report the results via Fig 5. We can see that joint training approach gives the upper bound of performance. CL approaches perform similarly. Further, GDumb strategy and naive training strategy gives lowest performance compared to others. Naive training strategy does not consider any measure to handle forgetting of past tasks and hence lower average performance. On the other hand, GDumb randomly consider a subset data from all encountered tasks for learning from scratch and hence perform poorly due to sample selection bias.
A detailed comparison of best performance achieved by various approaches is provided in Table III. Specifically, for each approach, we report forgetting measure by BWT, forward transfer by FWT, incremental learning capabilities by $A$, and final average accuracy by $ACC$. Here, also we can see that joint and cumulative approaches in non-CL category tend to outperform in all performance metrics as compared to CL approaches. However, assumption of availability of all the tasks together and retraining from scratch whenever a new tasks arrives, limits their applicability in real life. While comparing the CL approaches, we can see that LwF achieves better handles forgetting achieving the best BWT of -1.23. However, other metrics (FWT, $A$, and $ACC$) are lower compared to that achieved by Replay. This is due to the fact that knowledge distillation step in LwF limits the plasticity to learn new tasks. It usually performs well in situation involving less drift/ new data over the episodes (also evident from compromised results in CI scenario). Remarkably, replay outperforms other CL approach including LwF by achieving (FWT, $A$, and $ACC$) as $\{68.98,70.12,71.48\}$. Further, we can see that GDumb with same memory size as with Replay performs poor. GEM and A-GEM approaches are highly compute intensive compared to Replay and also achieve lower performance compared to replay. Although, these CL approaches do not show a contrasting superiority over another for DI scenario. Therefore, CL approaches may be picked based on application requirement, for example, a system demanding a solution with low memory will favour memory-free approaches such as EWC, LwF, or SI and the other way can favour Replay approach as it offers best overall performance.
@@ VI-C1 Performance analysis for DI scenario
TABLE IV: Performance Comparison for CI scenario. First / second best performance in CL categories indicated in red / blue , respectively. Bold: Upper bound.
@@
VI-C2 Performance analysis for CI scenario
Figure 6 shows the performance of the different approaches in terms of average accuracy on only seen classes after learning each task in CIS setting. It is expected that over the session performance would drop as more and more classes are added, making the learning difficult. However, Replay is able to maintain the performance to a steady level over the training session. We can see that it outperforms all other CL approaches, even surpassing non-CL strategies supposed to give upper-bound of performance like cumulative and joint.
We report detailed comparison of all approaches in terms of BWT, FWT, $A$, and $ACC$ in Table IV. From the table, we can observe that regularization-based approaches (EWC, LwF, SI) perform similar to the lower-bound approaches i.e., naive approach. We can observe that Replay approach achieves the best values of (BWT, FWT, $A$, $ACC$) as $\{-1.04,96.97,96.98,95.74\}$. It outperforms with an incremental learning accuracy (A) of 96.98% compared to other rehearsal approaches like GDumb, GEM, and A-GEM which achieve only 77.18%, 42.19% and 64.82%, respectively. We can see that GDumb, which is not a carefully designed rehearsal-based approach, acheives second best results and shows superior performance to complex rehearsal-based approaches like GEM and A-GEM.
From tables III and IV, we conclude that the rehearsal-based strategies perform well in both DI and CI scenarios while regularization approaches show inadequate performance in case of CI scenarios. This is attributed to that fact that CI scenario is the most difficult scenario as new task consists of novel classes and hence it is very hard to learn new tasks along with remembering past just with regularization step in the loss function. Some samples from past classes are important to rejuvenate the past class information in the current model.
TABLE V: Performance Comparison with different hyperparameters for DI scenario.
@@ VI-C2 Performance analysis for CI scenario
TABLE VI: Comparison of approaches on different hyperparameters for CI scenario.
@@
VII Discussion
For both DI and CI scenarios, we have conducted various experiments with different techniques, with a wide range of hyperparameters.
From Table V, we observe that different CL strategies perform similarly well in DI scenario. Among the regularization approaches (EWC, LwF, and SI), EWC achieves better performance (BWT, FWT, $A$, $ACC$) as the $\lambda$ value increases, while SI performs better at lower $\lambda$ values. LwF shows an increase in performance (BWT, FWT, $A$, $ACC$) as the $\alpha$ value decreases. EWC outperforms the other regularization approaches in terms of overall accuracy ($A$=67.67%) as well as average accurcay ($ACC$=71.93%), while LwF excels in reducing forgetting (BWT=-1.23) and improving future learning (FWT=65.85).
For rehearsal-based approaches (GDumb, Replay, GEM, and A-GEM), we can see that we can improve performance at the cost of increase memory size. GEM and A-GEM achieve similar performance to regularization approaches (EWC, SI, and LwF) but better than GDumb. Since the compute requirements in GEM and A-GEM are very high, so regularization approaches can be favoured over these. GDumb with lower memory size (mem=1000) show very poor performance and improves if we increase memory size to 2000. It achieves $A$ as 57% and 61.60% with memory size of 1000 and 2000, respectively. In contrast, Replay achieves better results than GDumb even with low memory. It records $A$ as 69.36% and 70.12% for memory size of 1000 and 2000, respectively.
In CI scenario, as shown in Table VI, we observe that regularization-based approaches (EWC, SI, and LwF) fail to mitigate forgetting, for a wide range of hyperparameters as for EWC and SI ($\lambda$ in {1, 2, 3}) and for LwF ($\{\alpha=1,T=1\}$, $\{\alpha=2,T=2\}$, $\{\alpha=3,T=3\}$). Conversely, rehearsal techniques perform relatively well. The performance should increase with increase in memory size but it is sometimes violated in CI scenario by GDumb, GEM and A-GEM. CI scenario is very hard compared to DI and thus a carefully designed strategy succeeds to handle the forgetting. We can observe GDumb perform relatively well compared to GEM and A-GEM. Still, Replay outperforms these approaches by a large margin in terms of all four metrics. On a wide range of memory values, Replay approach consistently outperforms all other techniques used in the study. The best performance in terms of final accuracy $ACC$ by Replay is 95.74% where as that with GDumb, GEM, and A-GEM is 53.15%, 39.00%, and 54.45%, respectively. Further, performance by Replay approach shows a very small decrement even when the memory size is reduced from 2000 samples to 500 samples. A-GEM performs better than GEM in terms of both accuracy and training time when varying the number of samples per task (200, 300, 400). Additionally, the training time for Replay with a memory size of 2000 is significantly less than that of GEM and A-GEM with smaller memory sizes of 1200 (200*6) samples.
@@ VII Discussion
VIII Conclusion
We investigate CL approaches on DI and CI scenarios with 6 sequential tasks curated from a large range of DCASE datasets for machine sound monitoring. With the extensive comparative analysis performed, we observe that the performance of rehearsal-based strategies outperforms regularization-based CL approaches in CI scenario whereas both categories of approaches perform seemingly well in DI scenario. We achieve overall accuracy (A) with Replay approach as 70.12% for the DI scenario and 96.98% for the CI scenario. Overall, this comparative analysis emphasizes the importance of selecting appropriate CL strategies adapted to the specific requirements and complexities of audio data, ensuring optimal performance and robustness in real-world applications.
There are different unexplored real-world challenges associated with the use of CL in audio based application like machine monitoring or farm surveillance. In the future, we aim to investigate CL in farm monitoring which is prone to encounter more challenging and diverse domain shifts due to change is weather, season, and dynamic background environment. Further, in real-life monitoring investigating CL with blurred domain boundaries are yet to be explored.

### v2

@@ III-B Non-Continual learning baselines
III-B1 Joint-training
This is a popular approach in the non-CL category, where all the training episodes are jointly trained one time and achieve the highest average performance and hence also regarded as an upper bound. However, this method assumes the availability of all the episodes simultaneously, which is a bottleneck.
III-B2 Cumulative learning
This non-CL training strategy is another widely adopted approach offering upper-bound performance. Here, the training process is started from scratch each time a new episode is available. It considers the current and all the previous episodes together for retraining. However, previously learned knowledge is wasted due to retraining from scratch each time a new episode is encountered, leading to high computation requirements.
III-B3 Naive
This non-CL strategy achieves lower-bound performance. Here, a traditional transfer learning mechanism is followed. Whenever a new episode is available, the model adapts itself to this new data without following any mechanism to control forgetting past data.
@@ III-B3 Naive
TABLE III: Details of datasets
@@ IV-B Dataset for CI Scenario
Contrary to the domain-incremental dataset discussed in the earlier section, here, the end goal is machine-type classification. In the class-incremental learning scenario, there needs to be the inclusion of classification classes over time. Hence, we opt to classify machine types as the DCASE datasets offer a large number of machine types, which can then be sequentially learned by the CL model. To focus only on machine-type classification, either normal or abnormal data should be considered. Since the normal class data is available in large amounts, hence we consider the normal class samples of different machine sounds here. Specifically, we curate six episodes by utilizing the training part (it has only normal samples) of the development segment in task-2 of DCASE2020 and the source domain of the additional training segment in task-2 of DCASE2023 challenges. The choice of these datasets is such that they have disjoint machine classes. Specifically, we consider ToyCar and ToyConveyer machine types from DCASE2020 in the first episode, Valve, and Fan machine types in the second episode, and then Pump and Slider machine types in the third episode. Then, from the DCASE2023, two machine types, viz., Vacuum and ToyTank, are considered in the fourth episode, ToyNscale and ToyDrone are considered in the fifth episode, and finally, the last episode has the remaining classes from DCASE2023, including Bandsaw, Grinder, and Shaker. Inside every episode, the number of samples for each class is considered in balance amount, i.e., the machine types have 3000, 2804, 990, 990, and 990 samples in the first, third, fourth, fifth, and sixth episodes, respectively. Except for the second episode in which the two machine types, namely Valve and Fan, have 3000 and 2804 samples, respectively. A visual representation of the class-incremental dataset is depicted in Fig. 3. The detailed description of datasets of both scenarios is summarized in Table III.
V Experiments
V-A Experimental setup
@@
In all the approaches, we keep the base model as a pre-trained ResNet50 [62] model initialized with ImageNet weights. We used Adam as an optimizer. In order to select the best hyperparameters for the classification model, multiple experiments were performed with a cumulative approach as it gives the upper bound. Specifically, we search best values for epochs in {10, 15,…50}, batch sizes in {4,8,…32}, and learning rates in {$041e-$ 04, … $051e-$ 05}. Best performance is achieved with a batch size of 8, a learning rate of $041e-$ 04, and epochs of 30 for the CI scenario. For the DI scenario, the best performance is achieved with a batch size of 8, a learning rate of $031e-$ 03, and epochs of 50. To facilitate a common evaluation scheme, the selected hyperparameters are used in all the approaches. Further, the hyperparameters for CL approaches, such as $\lambda$, memory size, etc., were also experimentally selected from a range of values. Specifically, for DI experiment, we set $\lambda=0.5$ in EWC, $\alpha=2$, $T=2$} in LwF, and $\lambda=0.8$ in SI. In case of CI experiment, $\lambda=2$ was set in EWC, $\alpha=2$, $T=2$} kept for LwF, and $\lambda=2$ for SI. Further, for both DI and CI, the memory sizes were kept as 2000 samples for the GDumb and Replay approaches. Audio clips were trimmed to 10 seconds, and then spectrograms were computed for each sample.
@@
V-B Evaluation metrics
Here, we describe the evaluation procedure and metrics in CL. Each of the training sessions is a classification problem; therefore, we compute accuracy on the testing data of each task. After completing the training session of $t^{th}$ task using training data $Tr_{t}$, we compute accuracy values on testing datasets $Te_{1},Te_{2},\dots$ yielding $T$ accuracy values. Thus, after completing all the training sessions on a stream of $T$ tasks, we get a train-test matrix $R^{T\times T}$, where $R_{t,j}$ represents accuracy on $j^{th}$ task after completing the training on $t^{th}$ task. An example of a train-test matrix with 5 tasks is shown in Fig. 4.
Apart from getting average accuracy across tasks, other important evaluation measures in CL include forgetting and forward transfer, which can be computed from this $R$ matrix. Below, we describe metrics for forward transfer, forgetting, and model accuracy.
Backward Transfer (BWT): BWT, a forgetting measure, quantifies the amount of influence of the current task on the previously learned task [37]. It shows the improvement in the performance of a task after learning new tasks. Thus, a positive value of BWT signifies that learning the current task $t$ has caused improvement in the performance of an old task $k<t$. On the other hand, a negative value of BWT indicates that learning the current task $t$ has caused degradation in the performance of an old task. Thus, an algorithm with greater values of BWT is superior to the one with small BWT. Blue and yellow cell values in Fig. 4 are utilized to compute BWT using the equation [63] below.
@@ V-B Evaluation metrics
Forward Transfer (FWT): FWT signifies the amount of influence of the current task on upcoming tasks. It shows the zero-shot learning capabilities of the CL model. The higher the value of FWT, the better the model. Red cell values in Fig. 4 are utilized to compute FWT [63] measure with the following equation.
@@
TABLE IV: Performance comparison for DI scenario. First / second best performance in CL categories indicated in red / blue, respectively. Bold: Upper bound.
@@
V-C Results
V-C1 Performance analysis for DI scenario
For the DI scenario, we compute average accuracy over all tasks after completing a training session and report the results via Fig 5. We can see that the joint training approach gives the upper bound of performance. CL approaches perform similarly. Further, the GDumb strategy and naive training strategy give the lowest performance compared to others. Naive training strategy does not consider any measure to handle forgetting of past tasks and hence lower average performance. On the other hand, GDumb randomly considers a subset of data from all encountered tasks for learning from scratch and hence performs poorly due to sample selection bias.
A detailed comparison of the best performance achieved by various approaches is provided in Table IV. Specifically, for each approach, we report forgetting measures by BWT, forward transfer by FWT, incremental learning capabilities by $A$, and final average accuracy by $ACC$. Here, we can also see that joint and cumulative approaches in the non-CL category tend to outperform in all performance metrics as compared to CL approaches. However, the assumption of availability of all the tasks together and retraining from scratch whenever a new task arrives limits their applicability in real life. While comparing the CL approaches, we can see that LwF better handles forgetting achieving the best BWT of -1.23. However, other metrics (FWT, $A$, and $ACC$) are lower compared to that achieved by Replay. This is due to the fact that the knowledge distillation step in LwF limits the plasticity to learn new tasks. It usually performs well in situations involving less drift/ new data over the episodes (also evident from compromised results in CI scenario). Remarkably, replay outperforms other CL approaches, including LwF, by achieving (FWT, $A$, and $ACC$) as $\{68.98,70.12,71.48\}$. Further, we can see that GDumb, with the same memory size as Replay, performs poorly. GEM and A-GEM approaches are highly compute-intensive compared to Replay and achieve lower performance than replay. However, these CL approaches do not show a contrasting superiority over another DI scenario. Therefore, CL approaches may be picked based on the application requirements; for example, a system demanding a solution with low memory will favor memory-free approaches such as EWC, LwF, or SI. Otherwise, way can favor the Replay approach as it offers the best overall performance.
@@ V-C1 Performance analysis for DI scenario
TABLE V: Performance Comparison for CI scenario. First / second best performance in CL categories indicated in red / blue , respectively. Bold: Upper bound.
@@
V-C2 Performance analysis for CI scenario
Figure 6 shows the performance of the different approaches in terms of average accuracy on only seen classes after learning each task in the CIS setting. It is expected that over the session, performance would drop as more and more classes are added, making learning difficult. However, Replay is able to maintain the performance to a steady level over the training session. We can see that it outperforms all other CL approaches, even surpassing non-CL strategies that are supposed to give an upper-bound performance, such as cumulative and joint.
We report a detailed comparison of all approaches in terms of BWT, FWT, $A$, and $ACC$ in Table V. From the table, we can observe that regularization-based approaches (EWC, LwF, SI) perform similarly to the naive approach. We can observe that the Replay approach achieves the best values of (BWT, FWT, $A$, $ACC$) as $\{-1.04,96.97,96.98,95.74\}$. It outperforms with an incremental learning accuracy (A) of 96.98% compared to other rehearsal approaches like GDumb, GEM, and A-GEM, which achieve only 77.18%, 42.19%, and 64.82%, respectively. We can see that GDumb, which is not a carefully designed rehearsal-based approach, achieves the second-best results and shows superior performance to complex rehearsal-based approaches like GEM and A-GEM.
From tables IV and V, we conclude that the rehearsal-based strategies perform well in both DI and CI scenarios, while regularization approaches show inadequate performance in the case of CI scenarios. This is attributed to the fact that the CI scenario is the most difficult scenario as the new task consists of novel classes, and hence, it is very hard to learn new tasks along with remembering the past just with the regularization step in the loss function. Some samples from past classes are important to rejuvenate the information from past classes in the current model.
TABLE VI: Performance Comparison with different hyperparameters for DI scenario.
@@ V-C2 Performance analysis for CI scenario
TABLE VII: Comparison of approaches on different hyperparameters for CI scenario.
@@
VI Discussion
From Table VI, we observe that different CL strategies perform equally well in DI scenario. Among the regularization approaches (EWC, LwF, and SI), EWC achieves better performance (BWT, FWT, $A$, $ACC$) as the $\lambda$ value increases, while SI performs better at lower $\lambda$ values. LwF shows an increase in performance (BWT, FWT, $A$, $ACC$) as the $\alpha$ value decreases. EWC outperforms the other regularization approaches in terms of overall accuracy ($A$=67.67%) as well as average accuracy ($ACC$=71.93%), while LwF excels in reducing forgetting (BWT=-1.23) and improving future learning (FWT=65.85).
For rehearsal-based approaches (GDumb, Replay, GEM, and A-GEM), we can see that we can improve performance at the cost of increasing memory size. GEM and A-GEM achieve similar performance to regularization approaches (EWC, SI, and LwF) but better than GDumb. Since the compute requirements in GEM and A-GEM are very high, so regularization approaches can be favoured over these. GDumb with lower memory size (mem=1000) shows very poor performance and improves if we increase memory size to 2000. It achieves $A$ as 57% and 61.60% with memory sizes of 1000 and 2000, respectively. In contrast, Replay achieves better results than GDumb even with low memory. It records $A$ as 69.36% and 70.12% for memory sizes of 1000 and 2000, respectively.
In the CI scenario, as shown in Table VII, we observe that regularization-based approaches (EWC, SI, and LwF) fail to mitigate forgetting, for a wide range of hyperparameters as for EWC and SI ($\lambda$ in {1, 2, 3}) and for LwF ($\{\alpha=1,T=1\}$, $\{\alpha=2,T=2\}$, $\{\alpha=3,T=3\}$). Conversely, rehearsal techniques perform relatively well. The performance should increase with the increase in memory size, but it is sometimes violated in CI scenarios by GDumb, GEM, and A-GEM. CI scenario is very hard compared to DI, and thus, a carefully designed strategy succeeds in handling the forgetting. We can observe that GDumb performs relatively better compared to GEM and A-GEM. Still, Replay outperforms these approaches by a large margin in terms of all four metrics. On a wide range of memory values, the Replay approach consistently outperforms all other techniques used in the study. The best performance in terms of final accuracy $ACC$ by Replay is 95.74% whereas that with GDumb, GEM, and A-GEM is 53.15%, 39.00%, and 54.45%, respectively. Further, performance by the Replay approach shows a very small decrement even when the memory size is reduced from 2000 samples to 500 samples. A-GEM performs better than GEM in terms of both accuracy and training time when varying the number of samples per task (200, 300, 400). Additionally, the training time for Replay with a memory size of 2000 is significantly less than that of GEM and A-GEM with smaller memory sizes of 1200 (200*6) samples.
@@ VI Discussion
VII Conclusion
We investigate CL approaches on DI and CI scenarios with 6 sequential tasks curated from a large range of DCASE datasets for machine sound monitoring. With the extensive comparative analysis performed, we observe that the performance of rehearsal-based strategies outperforms regularization-based CL approaches in the CI scenario, whereas both categories of approaches perform seemingly well in the DI scenario. We achieve overall accuracy (A) with the Replay approach as 70.12% for the DI scenario and 96.98% for the CI scenario. Overall, this comparative analysis emphasizes the importance of selecting appropriate CL strategies adapted to the specific requirements and complexities of audio data, ensuring optimal performance and robustness in real-world applications.
There are different unexplored real-world challenges associated with the use of CL in audio-based applications like machine monitoring or farm surveillance. In the future, we aim to investigate CL in farm monitoring, which is prone to encounter more challenging and diverse domain shifts due to changes in weather, season, and dynamic background environment. Further, in real-life monitoring, investigating CL with blurred domain boundaries is yet to be explored.